gvps65:

Tomas wrote the equation y = 3x +. When Sandra wrote her equation, they discovered that her equation had all the same solutions as Tomas’s equation. Which equation could be Sandra’s?

Vocaloid:

Something got cut off after the plus sign But anyway if Sandra’s has the same solutions then Sandra’s equation must be some scalar multiple of Tomas’s equation Put simpler, there must be some number you can multiply by Tomas’s to get Sandra’s So look at the answer choices and see if any choices fit this description

Cupcake123456:

4 y = 12 x + 3 12 x − 4 y + 3 = 0 Explanation: An equation can be given in many forms and still mean the same. y = 3 x + 3 4 (known as the slope/intercept form.) Multiplied by 4 to remove the fraction gives: 4 y = 12 x + 3 → 12 x − 4 y = − 3 (standard form) 12 x − 4 y + 3 = 0 (general form) These are all in the simplest form, but we could also have infinitely variations of them. 4 y = 12 x + 3 could be written as: 8 y = 24 x + 6 12 y = 36 x + 9 , 20 y = 60 x + 15 etc